# Metric Poincaré inequalities for graphs

### v2

@@ Abstract.
This article considers embeddings of bounded degree graphs into general metric spaces. Our first main result is a metric analogue of Matoušek’s extrapolation that relates the Poincaré constants $\gamma(G,\varrho^{p})$ and $\gamma(G,\varrho^{q})$ for any pair of exponents $0<p,q<\infty$, any bounded degree expander graph $G$, and any metric space $\mathcal{M}=(M,\varrho)$. Our second main result provides a sharp estimate of the Poincaré constant $\gamma(G,\varrho)$ in terms of the cardinalities of the vertex set of $G$ and the target metric space $\mathcal{M}=(M,\varrho)$, in the setting of random graphs. Both theorems utilize a novel structural dichotomy for metric embeddings of graphs.
These results lead to the resolution of several problems within the theory of metric embeddings. Among the applications are estimates on the nonlinear spectral gap of metric snowflakes, optimal estimates on the minimum cardinality of (bi-Lipschitz) universal metric spaces for graphs, and sharp lower bounds on the bi-Lipschitz distortion of random regular graphs into arbitrary metric spaces of large cardinality.
@@ 1. Introduction
A fruitful research direction—with origins in metric geometry, functional analysis, and geometric group theory—is to replace the Euclidean norm in (1.2) with other norms and, even more generally, arbitrary metrics. The result is a so-called nonlinear (i.e., non-Euclidean) Poincaré inequality. Early appearances of such inequalities can be found in the works of Enflo [E76], Bourgain–Milman–Wolfson [BMW86], Gromov [Gr83, Gr03], Pisier [Pi86], Linial–London–Rabinovich [LLR95], and Matoušek [Ma97]. We refer the reader to the introduction of Mendel–Naor [MN14] and the Bourbaki lectures of Eskenazis [Es22] for a detailed overview of this rich area and its numerous remarkable applications in metric geometry and algorithm design.
@@
The purpose of this article is to prove general metric analogues of core pieces of the theory of nonlinear spectral gaps for normed spaces, and to present related applications. A principal reason for studying the fully general setting comes from the Ribe program within metric geometry, which aims to develop metric analogues of normed space invariants. Other concrete and important motivations come from theoretical computer science and geometric group theory, which frequently consider graph embeddings into metric spaces, such as groups or other graphs, without normed structure.
@@ 1.1. Matoušek’s extrapolation for metric spaces
Despite the successful history of extrapolation principles for normed spaces, extending these results for general metric spaces has remained elusive. Analogues of Matoušek’s extrapolation within the specialized setting of Hadamard spaces have been recently studied by Eskenazis–Mendel–Naor [EMN25a], where they explicitly ask [EMN25a, Question 1.18] whether a version of Matoušek’s extrapolation holds under curvature assumptions. Our first main result resolves the extrapolation problem for all bounded degree expander graphs and all metric spaces.
@@ 1.2. Metric Poincaré inequalities for random graphs
A principle difficulty in estimating nonlinear spectral gaps for metric spaces is that many methods available for normed spaces are no longer applicable when the target space $\mathcal{M}$ is a general metric space. Thus, one either imposes extra structural assumptions on $\mathcal{M}$—which dramatically reduce the class of admissible metric spaces—or considers the statistical version of the problem where the domain is a random graph. Aiming for this latter setting, we note that (1.6) is also the strongest known general estimate even when the graph $G$ is random.
Our second main contribution is a sharp and purely metric analogue of the estimate (1.8) in the random graph setting. For the remainder of this paper, for any pair $n\geqslant d\geqslant 3$ of integers, $G(n,d)$ denotes the set of all $d$-regular graphs on $n]:=\{1,\dots,n\}$, and $\mathbb{P}_{G(n,d)}$ denotes the uniform probability measure on $G(n,d)$. We will always assume $dn$ is even.
@@
Remark 1.6.
The probabilistic estimate (1.9) is optimal for all values of the parameters $n,N$. This is discussed in Section 11.
Remark 1.7.
There is no deterministic analogue of Theorem 1.5 for general metric spaces. Indeed, by taking $\mathcal{M}$ equal to a deterministic bounded degree connected graph [^3] Graphs are viewed as metric spaces equipped with the shortest-path distance; see Subsection 3.1 for more details. $G$, we obtain that $\gamma(G,{\rm dist}_{G})\gtrsim\log\big(|V_{G}|\big)$, i.e., no improvement over (1.6) is available.
Remark 1.8 (Comparison of normed and metric settings).
Restricted to the setting of random graphs, our estimate for metric spaces (1.9) quantitatively recovers the optimal estimate for normed spaces (1.8). Indeed, assume that $X$ is a normed space with $\mathrm{dim}(X)\gtrsim\log n$. For a sufficiently large constant $C>0$, let $\mathcal{M}=(M,\|\cdot\|_{X})$ be an $n^{-C}$-net in the unit ball of $X$ with cardinality $|M|\leqslant(5\,n^{C})^{\mathrm{dim}(X)}$. By (1.9), with high probability, every map $f\colon[n]\to M$ satisfies (1.3) with a constant $\gamma=O\big(\min\big\{\log n,\log\log\big(|M|\big)\big\}\big)=O\big(\min\big\{\log n,\log\big(\mathrm{dim}(X)\big)\big\}\big)$. It is then easy to verify, by a standard approximation argument, that (1.3) also holds for all maps $f\colon[n]\to X$, implying (1.8).
1.3. Applications to metric embeddings
We proceed to present applications of our main results to metric embeddings.
@@ 1.3.1. Embeddings of random regular graphs into metric spaces
1.3.2. Bi-Lipschitz universality
The classical Fréchet embedding asserts that every $n$-point metric space embeds isometrically into $\ell_{\infty}^{k}$ for $k=n$. A fundamental problem originating in the seminal work of Johnson–Lindenstrauss [JL84] asks whether the dependence of $k$ on $n$ can be improved if we allow the embedding to have bi-Lipchitz distortion at most $D$. For brevity, such embeddings are said to be $D$-embeddings.
Upper bounds were obtained by Johnson–Lindenstrauss–Schechtman [JLS87] and, in a stronger form, by Matoušek [Ma92] who proved that, as long as $1\leqslant D\leqslant\frac{\log n}{\log\log n}$, any $n$-point metric space $D$-embeds into $\ell^{k}_{\infty}$ with $k\lesssim n^{C/D}$, where $C>0$ is a universal constant. A subsequent breakthrough of Matoušek [Ma96] confirmed that this is optimal via the following rigidity result for embeddings. For every $n$, there exists an $n$-point metric space $\mathcal{M}$, constructed as a regular graph of large girth and small diameter, such that if $\mathcal{M}$ admits a $D$-embedding into a normed space $X$, then necessarily $\dim(X)\gtrsim n^{c/D}$, where $c>0$ is a universal constant. An alternative and more robust proof was given by Naor [Na21], who used (1.7) to show that spectral expanders have the aforementioned rigidity. We refer the reader to [Na18] for a detailed overview, and to [AT25] for very recent progress in the regime $\Omega(\log n)\leqslant D\leqslant n$.
@@ 1.3.2. Bi-Lipschitz universality
Remark 1.14 (Universal graphs).
An interesting, analogous research direction in combinatorics seeks to identify the minimum cardinality of a “universal graph”. Given parameters $n$ (and $d$), a universal graph contains all graphs (of degree at most $d$) on $n$ vertices as induced subgraphs. We refer the interested reader to the recent sharp answers for this problem in the case of bounded degree $d$ and arbitrary degree, respectively, due to Alon–Nenadov [AN19] and Alon [Al17].
1.3.3. Metric snowflakes
We conclude with an application of our extrapolation result to the study of metric snowflakes. The $\epsilon$-snowflake of a metric $\mathcal{M}=(M,\varrho)$ is given by $\mathcal{M}_{\epsilon}=(M,\varrho^{1-\epsilon})$, where $\epsilon\in(0,1)$. The snowflake of a space often has better embeddability properties than the original space. The canonical example of this is the seminal embedding theorem of Assouad [As83], which roughly asserts that snowflakes of doubling metric spaces admit low-distortion embeddings into finite dimensional Euclidean space. Following Assouad’s theorem, snowflakes have become an important construction in metric geometry, acting as a metric notion of regularization; we refer to the survey [Es22] for a comprehensive discussion of applications.
Corollary 1.15 (Extrapolation for snowflakes).
In particular, for any fixed $\epsilon\in(0,1)$, a bounded degree expander graph $G$ satisfies a dimension-free metric Poincaré inequality with respect to $\mathcal{M}$ if and only if $G$ satisfies a dimension-free inequality for the $\epsilon$-snowflake of $\mathcal{M}$.
@@ 2.1. A structural dichotomy
A core observation enabling both our extrapolation and metric Poincaré results is the following structural dichotomy for functions $f\colon[n]\to M$. Such functions will be classified as either concentrated or non-concentrated. Non-concentrated functions turn out to always satisfy $\gamma(G,\varrho^{q};f)=O(1)$, for any $q\geqslant 1$ and any bounded degree expander graph $G$. On the other hand, concentrated functions exhibit structure that is amenable to analysis.
@@ 2.2. Extrapolation
We now sketch the proof of our extrapolation result, Theorem 1.2 (the full proof is given in Sections 45). Recall that our goal is to compare $\gamma(G,\varrho^{p})$ and $\gamma(G,\varrho^{q})$, where $G$ is a bounded degree expander graph and $0<p\leqslant q<\infty$. A simple argument enables us to restrict to $p\geqslant 1$.
@@ 2.2.2. Non-concentrated $f$
We now sketch the estimate $\gamma(G,\varrho^{q};f)=O(1)$ for any $q\geqslant 1$, any non-concentrated $f$, and any bounded degree expander graph $G$. A quantitative formulation is given in Proposition 4.3. We emphasize again that this is not a comparison inequality; it is a stronger, unconditional result.
@@ 12.2. Main result
such that for any connected $G\in G(n,d)$ with ${\rm diam}(G)\leqslant 500\log n$, we have $c_{\mathcal{M}}(G)\leqslant D$. Consequently, since $\min\big\{\log n,\log\log\big(|M|\big)\big\}=\Theta\big(\min\{\log n,\log\log k\}\big)$, by Corollaries 1.9 and 3.2, there exists $\tau^{\prime}=\tau^{\prime}(d)>0$ so that

### v3

@@ Abstract.
This article obtains purely metric counterparts of cornerstone results in the theory of embedding graphs into normed spaces. Our first main result is a metric analogue of Matoušek’s extrapolation relating the Poincaré constants $\gamma(G,\varrho^{p})$ and $\gamma(G,\varrho^{q})$ for any exponents $0<p,q<\infty$, any bounded-degree expander graph $G$, and any target metric space $\mathcal{M}=(M,\varrho)$. Our second main result provides a sharp estimate of the Poincaré constant $\gamma(G,\varrho)$ in terms of the cardinalities of the vertex set of $G$ and the metric space $\mathcal{M}=(M,\varrho)$, in the setting of random graphs. This yields optimal estimates on the minimum cardinality of (bi-Lipschitz) universal metric spaces for graphs, finally establishing a nonlinear analogue of Matoušek’s celebrated “incompressibility” theorem (1996). Further, we obtain estimates on the nonlinear spectral gap of metric snowflakes and sharp lower bounds on the distortion of random regular graphs into arbitrary metric spaces. Our proofs develop new nonlinear techniques, including random compression methods and a novel structural dichotomy for metric embeddings.
@@ 1. Introduction
A fruitful research direction—with origins in metric geometry, functional analysis, and geometric group theory—is to replace the Euclidean norm in (1.2) with other norms and, even more generally, arbitrary metrics. The result is a so-called nonlinear Poincaré inequality which serves as the main tool for establishing nonembedding results in metric geometry and geometric group theory. Early appearances of such inequalities can be found in the works of Enflo [E76], Bourgain–Milman–Wolfson [BMW86], Gromov [Gr83, Gr03], Pisier [Pi86], Linial–London–Rabinovich [LLR95], and Matoušek [Ma97]. We refer the reader to the introduction of Mendel–Naor [MN14] and the Bourbaki lectures of Eskenazis [Es22] for a detailed overview of this rich area and its numerous remarkable applications in metric geometry and algorithm design.
@@
The purpose of this article is to obtain purely metric counterparts to core pieces of the theory of spectral gaps for normed spaces, and to present related applications. A principal reason for studying the fully general setting comes from the Ribe program within metric geometry, which aims to develop metric analogues of normed space invariants. Other concrete and important motivations come from theoretical computer science and geometric group theory, which frequently consider graph embeddings into metric spaces, such as groups or other graphs, without normed structure. Among the applications obtained in this work are optimal estimates on the cardinality of universal metric spaces for graphs embeddings, which provide a natural counterpart to both classical and recent universality results across multiple branches of mathematics.
@@ 1.1. Matoušek’s extrapolation for metric spaces
Despite the successful history of extrapolation principles for normed spaces, extending these results for general metric spaces has remained elusive. Analogues of Matoušek’s extrapolation within the specialized setting of Hadamard spaces have been recently studied by Eskenazis–Mendel–Naor [EMN25a], where they explicitly ask [EMN25a, Question 1.18] whether a version of Matoušek’s extrapolation holds under curvature assumptions. Our first main result resolves the extrapolation problem for all bounded-degree expander graphs and all metric spaces.
@@ 1.2. Metric Poincaré inequalities for random graphs
The extent to which (1.8) is valid for general target metric spaces is a fundamental open problem in metric geometry, and in particular within the Ribe program [Na18]. A principal difficulty is that preexisting methods operate under strong assumptions on the target space, such as curvature or linear space structure. A second serious obstruction is that a naive nonlinear analogue of (1.8) cannot hold because $\mathcal{M}$ can be chosen adversarially depending on $G$; e.g., by taking $\mathcal{M}$ equal to a deterministic bounded-degree connected graph [^3] Graphs are viewed as metric spaces equipped with the shortest-path distance; see Subsection 3.1 for more details. $G$ we obtain that $\gamma(G,{\rm dist}_{G})\gtrsim\log\big(|V_{G}|\big)$, i.e., no improvement over (1.6) is available.
Our second main contribution is the following sharp asymptotic characterization of nonlinear spectral gaps for general metric spaces when $G$ is random. In particular, the second mentioned obstacle turns out to be the only obstruction to efficiently controlling the nonlinear Poincaré constant of graph embeddings. For the remainder of this paper, for any pair $n\geqslant d\geqslant 3$ of integers, $G(n,d)$ denotes the set of all $d$-regular graphs on $n]:=\{1,\dots,n\}$, and $\mathbb{P}_{G(n,d)}$ denotes the uniform probability measure on $G(n,d)$. We will always assume $dn$ is even.
@@
The proof of this theorem requires developing new methods suitable for the setting of general metric spaces; an extensive discussion is provided in Section 2.
Remark 1.6 (Optimality).
The estimate on the Poincaré constant in (1.9) is optimal for all values of the parameters $n,N$. This is discussed in Section 11. We also note that the probabilistic tail estimate in (1.9) is independent of $N$.
Remark 1.7 (Normed space setting, revisited).
Theorem 1.5 recovers the optimal estimate (1.8) for normed spaces when restricted to the setting of random bounded-degree graphs, despite being a fully metric result. Indeed, assume that $X$ is a normed space with $\mathrm{dim}(X)\gtrsim\log n$. For a sufficiently large constant $C>0$, let $\mathcal{M}=(M,\|\cdot\|_{X})$ be an $n^{-C}$-net in the unit ball of $X$ with cardinality $|M|\leqslant(5\,n^{C})^{\mathrm{dim}(X)}$. By (1.9), with high probability, every map $f\colon[n]\to M$ satisfies (1.3) with a constant $\gamma=O\big(\min\big\{\log n,\log\log\big(|M|\big)\big\}\big)=O\big(\min\big\{\log n,\log\big(\mathrm{dim}(X)\big)\big\}\big)$. It is then easy to verify, by a standard approximation argument, that (1.3) also holds for all maps $f\colon[n]\to X$, implying (1.8).
1.3. Applications to metric embeddings: bi-Lipschitz distortion
Our first main application is an estimate of the bi-Lipschitz distortion of embedding random graphs into arbitrary target metric spaces, yielding an optimal characterization for a wide range of parameters. Estimates on the spectral gap of metric snowflakes are also given.
@@ 1.3.1. Embeddings of random regular graphs into metric spaces
1.3.2. Metric snowflakes
Next, we consider an application of our extrapolation result to the study of metric snowflakes. The $\epsilon$-snowflake of a metric $\mathcal{M}=(M,\varrho)$ is given by $\mathcal{M}_{\epsilon}=(M,\varrho^{1-\epsilon})$, where $\epsilon\in(0,1)$. The snowflake of a space often has better embeddability properties than the original space. The canonical example of this is the seminal embedding theorem of Assouad [As83], which roughly asserts that snowflakes of doubling metric spaces admit low-distortion embeddings into finite dimensional Euclidean space. Following Assouad’s theorem, snowflakes have become an important construction in metric geometry, acting as a metric notion of regularization; we refer to the survey [Es22] for a comprehensive discussion of applications.
Corollary 1.11 (Extrapolation for snowflakes).
In particular, for any fixed $\epsilon\in(0,1)$, a bounded-degree expander graph $G$ satisfies a dimension-free metric Poincaré inequality with respect to $\mathcal{M}$ if and only if $G$ satisfies a dimension-free inequality for the $\epsilon$-snowflake of $\mathcal{M}$.
1.4. Applications to metric embeddings: universal spaces
Our second main application is a sharp estimate on the cardinalityof (bi-Lipschitz) universal metric spaces, establishing a purely metric counterpart to Matoušek’s celebrated “incompressibility” theorem [Ma96].
The search for universal structures (spaces containing all substructures from a given class) is a fundamental problem within many branches of mathematics. Classical examples include the Whitney and Nash embedding theorems in differential geometry and the Menger–Nöbeling theorem in topology. In combinatorics, recent breakthroughs of Alon–Nenadov [AN19] and Alon [Al17] identify the minimum cardinality universal graphs which contain all subgraphs of a given size.
In the setting of metric embeddings, the classical Fréchet theorem asserts that every $n$-point metric space embeds isometrically into $\ell_{\infty}^{k}$ for k=n1k=n-1. A fundamental problem originating in the seminal work of Johnson–Lindenstrauss [JL84] asks about the minimum dimension universal space with respect to embeddings with bi-Lipchitz distortion at most $D$. For brevity, such embeddings are said to be $D$-embeddings.
Upper bounds were obtained by Johnson–Lindenstrauss–Schechtman [JLS87] and, in a stronger form, by Matoušek [Ma92] who proved that, as long as $1\leqslant D\leqslant\frac{\log n}{\log\log n}$, any $n$-point metric space $D$-embeds into $\ell^{k}_{\infty}$ with $k\lesssim n^{C/D}$, where $C>0$ is a universal constant. Matoušek’s celebrated “incompressibility” theorem [Ma96, Ma02] yields a matching lower bound, obtained via the following rigidity result for embeddings. For every $n$, there exists an $n$-point metric space $\mathcal{M}$, constructed as a regular graph of large girth and small diameter, such that if $\mathcal{M}$ admits a $D$-embedding into a normed space $X$, then necessarily $\dim(X)\gtrsim n^{c/D}$, where $c>0$ is a universal constant. An alternative and more robust proof was given by Naor [Na21], who used (1.7) to show that spectral expanders have the aforementioned rigidity. We refer the reader to [Na18] for a detailed overview, and to [AT25] for very recent progress in the regime $\Omega(\log n)\leqslant D\leqslant n$.
@@ 2.1. A structural dichotomy
A core observation enabling both our extrapolation and metric Poincaré results is the following structural dichotomy for functions $f\colon[n]\to M$. Such functions will be classified as either concentrated or non-concentrated. Non-concentrated functions turn out to always satisfy $\gamma(G,\varrho^{q};f)=O(1)$, for any $q\geqslant 1$ and any bounded-degree expander graph $G$. On the other hand, concentrated functions exhibit structure that is amenable to analysis.
@@ 2.2. Extrapolation
We now sketch the proof of our extrapolation result, Theorem 1.2 (the full proof is given in Sections 45). Recall that our goal is to compare $\gamma(G,\varrho^{p})$ and $\gamma(G,\varrho^{q})$, where $G$ is a bounded-degree expander graph and $0<p\leqslant q<\infty$. A simple argument enables us to restrict to $p\geqslant 1$.
@@ 2.2.2. Non-concentrated $f$
We now sketch the estimate $\gamma(G,\varrho^{q};f)=O(1)$ for any $q\geqslant 1$, any non-concentrated $f$, and any bounded-degree expander graph $G$. A quantitative formulation is given in Proposition 4.3. We emphasize again that this is not a comparison inequality; it is a stronger, unconditional result.
@@ 12.2. Main result
such that for any connected $G\in G(n,d)$ with ${\rm diam}(G)\leqslant 500\log n$, we have $c_{\mathcal{M}}(G)\leqslant D$. Consequently, since $\min\big\{\log n,\log\log\big(|M|\big)\big\}=\Theta\big(\min\{\log n,\log\log k\}\big)$, by Corollaries 1.8 and 3.2, there exists $\tau^{\prime}=\tau^{\prime}(d)>0$ so that